Daniel is 28 years younger than Kevin. Kevin and Daniel first met 3 years ago. Eight years ago, Kevin was 5 times as old as Daniel. How old is daniel now?
Answer by algebrahouse.com(1659)

k = Kevin's age now
k - 28 = Daniel's age now

The fact that they met 3 years ago means nothing.

k - 8 = Kevin's age 8 years ago
k - 36 = Daniel's age 8 years ago

k - 8 = 5(k - 36) {8 years ago, Kevin was 5 times as old as Daniel}
k - 8 = 5k - 180 {used distributive property}
-8 = 4k - 180 {subtracted k from each side}
4k = 172 {added 180 to each side}
k = 43 {divided each side by 4}
k - 28 = 15 {substituted 43, in for k, into k - 28}

Daniel is 15 now.

For more help from me, visit: www.algebrahouse.com